OpenStudy (anonymous):

An equation is shown below: 6(3x – 7) = 2 Which of the following correctly shows the beginning steps to solve this equation? (4 points) Step 1: 9x – 1 = 2 Step 2: 9x = 3 Step 1: 9x – 7 = 2 Step 2: 9x = 9 Step 1: 18x – 7 = 2 Step 2: 18x = 9 Step 1: 18x – 42 = 2 Step 2: 18x = 44

OpenStudy (anonymous):

\(\Huge\color{lightblue}{Welcome~To~Openstudy!!!}\)

OpenStudy (anonymous):

wich one is right the first second third or fourth

OpenStudy (anonymous):

first rellease the brackets 18x-42=2

OpenStudy (anonymous):

then add to both sides 42 18x-42+42=2+42 18x=44

OpenStudy (anonymous):

the last one is right

OpenStudy (anonymous):

so is it the first one?

OpenStudy (anonymous):

oh ok thanks i will medal you

OpenStudy (anonymous):

np :)